OpenStudy (anonymous):

Differentiate 3x-e^(x-2)-4

OpenStudy (lalaly):

\[\frac{d}{dx}(2x-e^{x-2}-4)=2-e^{x-2}\]

OpenStudy (lalaly):

sorry meant 3x

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

damn that's pro lalaly

OpenStudy (lalaly):

ur welcome

OpenStudy (anonymous):